OpenStudy (anonymous):

Use the quadratic formula to solve for 2x^2-7x-5=0

OpenStudy (accessdenied):

Well, what is the quadratic formula for a general quadratic: \(ax^2 + bx + c = 0\)?

OpenStudy (accessdenied):

Once we have that established, we simply have to plug in the corresponding a, b, and c into the formula and evaluate.

OpenStudy (accessdenied):

The quadratic formula is: \( \displaystyle x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a} \) From here, it is simply taking the a, b, and c from the original equation and inserting them. Please respond if you are having any issues. :)